# The Still Secret Ballot:

### 06 Mar 2024 v2

@@ Abstract
After an election, should officials release an electronic record of each ballot? The release of ballots could bolster the legitimacy of the result. But it may also facilitate vote revelation, where an analyst unravels the secret ballot by uniquely linking votes on an anonymous ballot to the voter’s name and address in the public voter file. We first provide a theoretical model of how vote revelation could occur under various election-reporting regimes. Perhaps counterintuitively, releasing ballot records is no more revelatory than the typical practice of releasing aggregate vote tallies. We then present the first empirical evaluation of vote revelation, using the 2020 election in Maricopa County, Arizona, as a case study. For 99.8% of voters, the release of ballot records led to no revelation of any vote choice. We conclude the ballot can be both public and still as secret as it is under the current reporting practices.
@@ Introduction
Third, and finally, the extent of revelation is directly relevant to current litigation about whether cast vote records should be public record for transparency or kept confidential because of privacy. The extent of revelation is also pertinent to policy decisions about how to best strike the difficult balance between transparency and privacy in election reporting. We estimate that less than 3,500 out of over 2 million Maricopa County voters could have any of their votes revealed. Further, when votes are revealed, they are disproportionately provisionally cast ballots, which only comprise 0.1 percent of total ballots cast, and are usually votes for candidates who are popular in the area in which the voter lives. These findings help inform whether a particular response—ranging from not releasing any individual ballots to aggregating, suppressing, or adding random noise to results—is proportionate to the potential privacy loss.
@@
To understand the debate about ballot records, it is helpful to conceive of election reporting as a query of an election database. The building blocks of an election database include a table of individual ballot records and a table of individual voter registration records. The ballot records indicate each choice made in each contest by each voter in any election, while the voter registration records enumerate which registrants cast a ballot in the same election. Importantly, the ballot records are anonymous, i.e., no names are attached to the actual ballot, but voter records are not. In this framework, the empirical concern raised by public officials is that a vote choice recorded in a table of election results will be linked to an individual record in a table of voters. We call this vote revelation.
@@
Our conception of vote revelation is meant to be responsive to the particular empirical concerns of public officials discussed above. If public officials fear that releasing individual ballot records will reveal vote choices, it is useful to understand the extent to which that concern is realized. However, our conception of vote revelation is neither a sufficient nor necessary condition for determining whether any method of election reporting violates the legal requirement for a secret ballot. It is not a necessary condition because, in general, there is no singular, well-defined secret ballot. Instead, the secret ballot is a state right, enshrined in state constitutions and state statutes in different ways. For example, some state constitutions guarantee “absolute” secrecy and some simply guarantee that voting “shall be by ballot” [25]. Similarly, some permit voters to waive the right and some do not [25]. Given the variation in text and other constitutional commitments, courts—even if not election officials—may also consider other dimensions of the secret ballot, beyond the extent of revelation, such as the harm of revelation or risk of revelation. Vote revelation is also not a sufficient condition for the secret ballot, for the simple reason that public officials still need to supply their own threshold to distinguish some vote revelation from too-much vote revelation. Nonetheless, we argue that our concept of vote revelation is an important yardstick by which to measure the secret ballot.
@@
We distinguish between two different types of vote revelation. A vote can either be revealed to (1) the public as a whole or (2) to specific local citizens who are able to supplement the election results and the voter file with knowledge of their own vote (or neighbors’ votes). We term the former public vote revelation in contrast to the latter local vote revelation. Both types of revelation focus on a reporting unit for election results defined by some combination quasi-identifiers, either precinct, ballot style, or vote method.
@@
In formal privacy terminology, the concept of local revelation makes clear that datasets which are $\ell$-diverse are not necessarily immune from vote revelation [14]. Instead, a person who voted themselves can use their own private knowledge to back out how others voted. For this reason, the mechanism for what we call local vote revelation is known as a background knowledge attack [47].
@@
One concern is that ballots could reveal more information because they show how a single voter voted on all offices. But the example in Figure 2 already showed otherwise. Even if we had individual ballots for the 10 in-person voters, we could not identify how a particular individual voted on the tax referendum even though we could identify that they voted for Trump. In general, individual ballot records have been mischaracterized as posing an exceptional privacy threat. But releasing ballot records can reveal no more votes than would otherwise be revealed by the most granular level of aggregate reporting. [^3] Our claim is specific to our definition of vote revelation. Our definition of vote revelation does not include so-called negative revelation: the ruling out of choices that a voter may have made. Negative revelation is often found in granular election results where it is not uncommon to find zero votes cast for particular minor-party candidates in a reporting unit ruling out the possibility that anyone known to have voted in that reporting unit support that candidate. To be clear, cast vote records do facilitate more negative revelation than aggregate election results. Consider the same two contests as Figure 2, the President and a yes/no referendum. Suppose a vote buyer asks one voter to vote [Trump; Yes] and another to vote [Biden; No]. However, if the first voter voted [Trump; No], and the second voted [Biden; Yes], the cast vote records would reveal that both voters had reneged, while aggregate results would not. We thank Samuel Baltz for providing this example.
@@
From a privacy perspective, the relevant difference is that cast vote records only contain a voter’s contest choices, while a ballot image would capture all marks or notations made on the ballot, including those that do not indicate a choice for a particular candidate. For this reason, voters are often prohibited from adding identifying marks to their ballot to identify themselves (See, e.g., Minn. Stat. §204C.22(13)), and election officials attempt to redact ballot images with identifying marks [[]also see Figure B2 for a redacted mark]Colorado2018. Nonetheless, ballot images are more vulnerable to vote revelation than cast records because deterrence and redaction may be ineffective or incomplete. Further, as we describe below, our interest is not in self-revelation, but rather the revelation of vote choices that were not intended to be revealed to anyone. For this reason, we study what the choices themselves can reveal. Those choices are contained in the cast voter records and, for the remainder of this article, we focus specifically those records.
@@
Second, we do not consider the possibility of vote revelation from quasi-identifiers other than precinct, ballot style, and vote method. For example, election officials have voiced concern that including time-stamps on ballots records indicating exactly when each ballot was cast (or in what order the ballots were cast) could be used to reveal votes (See, e.g., Honey v. Lycoming County). [^5] In fact, at least one voting system is known to have inadvertently disclosed the order in which in-person ballots were cast [18]. Assuming the order in which individuals voted is observable or recorded in a poll book, the ability to sort the ballots in the order in which they were cast is fully revealing of which voters cast which ballot. We set aside such concerns arising from other quasi-identifiers because the solution to these concerns is about election administration, not the reporting of election results. To continue the example, election officials need not provide a time-stamp of a voter’s interaction with a ballot in a publicly available cast vote record. In contrast, if election officials are required to report precinct-level results, they must necessarily associate the voter’s precinct with each ballot in order to administer the election. Similarly, information about the ballot style is inherent to each ballot and cannot be redacted.666 The cast vote records produced by some voting systems do not connect vote choices made by voters across a multiple-page ballot. In this case, ballot style may not be entirely revealed by the cast vote records and, thus, ballot style information is not inherent to cast vote records in those cases. In the limit, a cast vote record could report vote choices by contest and not link voter choices across any pair of contests for any voter. Such a cast vote records database would be informationally equivalent to a database that contained only candidate vote totals. Election officials have also voiced concern about ballots in different languages, but we are not presently aware of ballot language being recorded in a voter file.
@@
Precinct 743 33 602 1,090 2,600
@@
Ballot Style & 381 1 7 41 1,897 5,484.4 Precinct $\times$ Vote Method 2,195 1 3 5 208 952.0 Ballot equivalent 4,397 1 1 1 11 475.2     … only mail method 1,742 1 2 4 257 1,099.6     … only in-person method 1,389 1 1 1 87 120.9     … only provisional method 1,320 1 1 1 3 4.7 Cast vote record 2,089,563 We use Ballot to refer to Precinct $\times$ Ballot Style $\times$ Vote Method units, as we describe in the main text.

### 09 Apr 2024 v3

@@ Abstract
After an election, should officials release an electronic record of each ballot? The release of ballots could bolster the legitimacy of the result. But it may also facilitate vote revelation, where an analyst unravels the secret ballot by uniquely linking votes on an anonymous ballot to the voter’s name and address in the public voter file. We first provide a theoretical model of how vote revelation could occur under various election-reporting regimes. Perhaps counterintuitively, releasing ballot records is no more revelatory than the typical practice of releasing aggregate vote tallies by precinct and method. We then present the first empirical evaluation of vote revelation, using the 2020 election in Maricopa County, Arizona, as a case study. For 99.8% of voters, the release of ballot records led to no revelation of any vote choice. We conclude the ballot can be both public and still as secret as it is under typical reporting practices.
@@ Introduction
Third, and finally, the extent of revelation is directly relevant to current litigation about whether cast vote records should be public record for transparency or kept confidential for privacy reasons. The extent of revelation is also pertinent to policy decisions about how to best strike the difficult balance between transparency and privacy in election reporting. We estimate that less than 3,500 out of over 2 million Maricopa County voters could have any of their votes revealed. Further, when votes are revealed, they are disproportionately provisionally cast ballots ( which only comprise 0.1 percent of total ballots cast), and are usually votes for candidates who are popular in the area in which the voter lives. These findings help inform whether a particular response—ranging from not releasing any individual ballots to aggregating, suppressing, or adding random noise to results—is proportionate to the potential privacy loss.
@@
To understand the debate about ballot records, it is helpful to conceive of election reporting as a query of an election database. The building blocks of an election database include a table of individual ballot records and a table of individual voter registration records. The ballot records indicate each choice made in each contest by each voter in any election, while the voter registration records enumerate which registrants cast a ballot in the same election. Importantly, the ballot records are anonymous, i.e., no names are attached to the actual ballot222 One caveat is that North Carolina marks some ballots—those cast during early voting—with an identifier such that it is “retrievable” (N.C. Gen. Stat. §163-166.45). To be clear, “[t]he ballot number or equivalent identifier shall not be a public record.” Instead, the ballot identifier allows election officials to remove a ballot from counting if it is later found to be fraudulent. While making the ballot retrievable means that it may no longer be anonymous to election officials, it is a criminal offense for anyone “who has access to an official voted ballot or record and knowingly discloses … how an individual has voted that ballot” (N.C. Gen. Stat. §163-165.1(e))., but voter records are not. In this framework, the empirical concern raised by public officials is that a vote choice recorded in a table of election results will be linked to an individual record in a table of voters. We call this vote revelation.
@@
Our conception of vote revelation is meant to be responsive to the particular empirical concerns of public officials discussed above. If public officials fear that releasing individual ballot records will reveal vote choices, it is useful to understand the extent to which that concern is realized. However, our conception of vote revelation is neither a sufficient nor necessary condition for determining whether any method of election reporting violates the legal requirement for a secret ballot. It is not a necessary condition because, in general, there is no singular, well-defined secret ballot. Instead, the secret ballot is a state right, enshrined in state constitutions and state statutes in different ways. For example, some state constitutions guarantee “absolute” secrecy and some simply guarantee that voting “shall be by ballot” [25]. Similarly, some permit voters to waive the right and some do not [25]. Given the variation in text and other constitutional commitments, courts—even if not election officials—may also consider other dimensions of the secret ballot, beyond the extent of revelation, such as the harm of revelation or risk of revelation. Vote revelation is also not a sufficient condition for the secret ballot, because public officials still need to supply their own threshold to distinguish some vote revelation from too much vote revelation. Nonetheless, we argue that our concept of vote revelation is an important yardstick by which to measure the secret ballot.
@@
We distinguish between two different types of vote revelation. A vote can either be revealed to (1) the public as a whole or (2) to specific local citizens who are able to supplement the election results and the voter file with knowledge of their own vote (or neighbors’ votes). We term the former public vote revelation in contrast to the latter local vote revelation. Both types of revelation focus on a reporting unit for election results defined by some combination of quasi-identifiers, either precinct, ballot style, or vote method.
@@
In formal privacy terminology, the concept of local revelation makes clear that datasets that are $\ell$-diverse are not necessarily immune from vote revelation [14]. Instead, a person who voted themselves can use their own private knowledge to back out how others voted. For this reason, the mechanism for what we call local vote revelation is known as a background knowledge attack [47].
@@
One concern is that ballots could reveal more information because they show how a single voter voted on all offices. But the example in Figure 2 already showed otherwise. Even if we had individual ballots for the 10 in-person voters, we could not identify how a particular individual voted on the tax referendum even though we could identify that they voted for Trump. In general, individual ballot records have been mischaracterized as posing an exceptional privacy threat. But releasing ballot records can reveal no more votes than would otherwise be revealed by the most granular level of aggregate reporting. [^4] Our claim is specific to our definition of vote revelation. Our definition of vote revelation does not include so-called negative revelation: the ruling out of choices that a voter may have made. Negative revelation is often found in granular election results where it is not uncommon to find zero votes cast for particular minor-party candidates in a reporting unit, ruling out the possibility that anyone known to have voted in that reporting unit supports that candidate. To be clear, cast vote records do facilitate more negative revelation than aggregate election results. Consider the same two contests as Figure 2, the President and a yes/no referendum. Suppose a vote buyer asks one voter to vote [Trump; Yes] and another to vote [Biden; No]. However, if the first voter voted [Trump; No], and the second voted [Biden; Yes], the cast vote records would reveal that both voters had reneged, while aggregate results would not. We thank Samuel Baltz for providing this example.
@@
From a privacy perspective, the relevant difference is that cast vote records only contain a voter’s contest choices, while a ballot image would capture all marks or notations made on the ballot, including those that do not indicate a choice for a particular candidate. For this reason, voters are often prohibited from adding identifying marks to their ballot to identify themselves (See, e.g., Minn. Stat. §204C.22(13)), and election officials attempt to redact ballot images with identifying marks [[]also see Figure B2 for a redacted mark]Colorado2018. Nonetheless, ballot images are more vulnerable to vote revelation than cast records because deterrence and redaction may be ineffective or incomplete. Further, as we describe below, our interest is not in self-revelation, but rather the revelation of vote choices that were not intended to be revealed to anyone. For this reason, we study what the choices themselves can reveal. Those choices are contained in the cast voter records and, for the remainder of this article, we focus specifically on those records.
@@
Second, we do not consider the possibility of vote revelation from quasi-identifiers other than precinct, ballot style, and vote method. For example, election officials have voiced concern that including time-stamps on ballots records indicating exactly when each ballot was cast (or in what order the ballots were cast) could be used to reveal votes (See, e.g., Honey v. Lycoming County). [^6] In fact, at least one voting system is known to have inadvertently disclosed the order in which in-person ballots were cast [18]. Assuming the order in which individuals voted is observable or recorded in a poll book, the ability to sort the ballots in the order in which they were cast is fully revealing of which voters cast which ballot. We set aside such concerns arising from other quasi-identifiers because the solution to these concerns is about election administration, not the reporting of election results. To continue the example, election officials need not provide a time-stamp of a voter’s interaction with a ballot in a publicly available cast vote record. In contrast, if election officials are required to report precinct-level results, they must necessarily associate the voter’s precinct with each ballot in order to administer the election. Similarly, information about the ballot style is inherent to each ballot and cannot be redacted.777 The cast vote records produced by some voting systems do not connect vote choices made by voters across a multiple-page ballot. In this case, ballot style may not be entirely revealed by the cast vote records and, thus, ballot style information is not inherent to cast vote records in those cases. In the limit, a cast vote record could report vote choices by contest and not link voter choices across any pair of contests for any voter. Such a cast vote records database would be informationally equivalent to a database that contained only candidate vote totals. Election officials have also voiced concern about ballots in different languages, but we are not presently aware of ballot language being recorded in a voter file.
@@
Precinct 743 33 602 1,090 2,600 2,812.3
Ballot Style 381 1 7 41 1,897 5,484.4
Precinct $\times$ Vote Method 2,195 1 3 5 208 952.0
Ballot equivalent 4,397 1 1 1 11 475.2
    … only mail method 1,742 1 2 4 257 1,099.6
    … only in-person method 1,389 1 1 1 87 120.9
    … only provisional method 1,320 1 1 1 3 4.7
Cast vote record 2,089,563
@@
We use Ballot to refer to Precinct $\times$ Ballot Style $\times$ Vote Method units, as we describe in the main text.